# WALT3D: Generating Realistic Training Data from Time-Lapse Imagery for Reconstructing Dynamic Objects under Occlusion

### v1

@@ Abstract
denotes equal contribution
@@ 2 Related Works
3D Reconstruction Under Occlusion: Reconstruction under severe occlusion is still in the nascent stages of research. Most algorithms developed focus on self-occluded objects with shape completion from partial observations [80, 11, 54]. On the other hand, shape models fitting for objects only with the visible regions either from images [44, 61, 32, 27] or depth sensors [1, 55, 73] have been explored. Due to the inherent challenges in annotating 3D information, ground-truth data for 3D object understanding is limited, often confined to specific object sets [42, 71], indoor [5], or driving scenarios [22, 6, 64], and struggles to generalize well to novel viewpoints such as stationary traffic cameras. Recent methods addressing pose estimation for vehicles [52, 70] and people [25, 39] have demonstrated robustness in handling occlusion, either implicitly or explicitly. Thus, our dataset generation method can significantly enhance the robustness of these approaches to occlusion.
@@ 3 Generating Realistic Supervision Data
In the context of human reconstruction, we employ the Human Mesh Recovery (HMR) method HMR 2.0 from Goel et al. [25]. Using HMR 2.0, we predict the SMPL pose and shape parameters for each detected unoccluded human [51]. To determine camera translation, we find the intersection between the backprojected ray from the bottom of the 2D bounding box with the ground plane. Given the camera’s intrinsic matrix $K$, ground plane equation $\begin{bmatrix}\mathbf{n}&d\end{bmatrix}$, and $p_{b}$ as the pixel coordinate of the bottom of the box, its depth is computed as $z_{b}=-\frac{d}{\mathbf{n}^{T}(K^{-1}p_{b})}$. For camera intrinsic parameters and ground plane equation, we employ a recent method by Vuong et al. [69] that utilizes Google Street View [26] to automatically calibrate stationary traffic cameras and infer the ground plane. The mined unoccluded object image with its corresponding segmentation mask, keypoint locations, and 3D poses are used in a clip-art based framework to generate 2D and 3D supervision data in severe occlusions as illustrated in Fig. 2.
@@ 4 Learning Reconstruction under Occlusion
3D Object Reconstruction: To regress object 3D pose and shape under occlusion, we extend the base network with a 3D regression branch. Each vehicle $N$ semantic 3D keypoints $\mathbf{X}$ are represented by a linear combination of the mean shape $\mathbf{\bar{Q}}$ and $K$ principal components $\mathbf{Q}_{1\dots K}$, where $\alpha_{k}$ is the shape coefficients: $\mathbf{X}=\mathbf{\bar{Q}}+\sum_{k=1}^{K}\alpha_{k}\mathbf{Q}_{k}$. Thus, we want to regress for its 6-DoF pose $\{(\mathbf{R},\mathbf{t})\in\mathbf{SE}(3)\}$ and shape coefficients $\alpha_{k}$. Given the predicted 2D keypoints and the corresponding 3D keypoints $\mathbf{X}$, we can obtain its 6-DoF pose through a differentiable PnP layer [9]. Defining $\pi(\cdot)$ as the projection function with known intrinsic $K$, we can optimize for the keypoint reprojection loss $\mathcal{L}_{\texttt{kp2D}}$, shape loss $\mathcal{L}_{\texttt{shape}}$, and pose loss $\mathcal{L}_{\texttt{pose}}$ as:
@@ 6.3 Ablation Analysis
Cross-evaluations on pedestrian tracking: We evaluate how our WALT3D data helps pedestrian tracking in Table 2. To better highlight how better detections lead to better tracking, we use a simple tracker SORT [4] with two detectors: one pretrained on COCO and another finetuned with our WALT3D data. On MOT17-train data, the model finetuned with our WALT3D data produces high-quality detections under strong occlusions, resulting in improvements over the baseline in all metrics, particularly favoring better detection quality like MOTA and DetA.

### v2

@@ Abstract
denotes equal contribution and joint first author
@@ 2 Related Works
3D Reconstruction Under Occlusion: Reconstruction under severe occlusion is still in the nascent stages of research. Most algorithms developed focus on self-occluded objects with shape completion from partial observations [85, 12, 57]. On the other hand, shape models fitting for objects only with the visible regions either from images [45, 65, 33, 62, 28] or depth sensors [1, 58, 11, 77] have been explored. Due to the inherent challenges in annotating 3D information, ground-truth data for 3D object understanding is limited, often confined to specific object sets [43, 75], indoor [5], or driving scenarios [23, 6, 68], and struggles to generalize well to novel viewpoints such as stationary traffic cameras. Recent methods addressing pose estimation for vehicles [54, 74] and people [26, 40] have demonstrated robustness in handling occlusion, either implicitly or explicitly. Thus, our dataset generation method can significantly enhance the robustness of these approaches to occlusion.
@@ 3 Generating Realistic Supervision Data
In the context of human reconstruction, we employ the Human Mesh Recovery (HMR) method HMR 2.0 from Goel et al. [26]. Using HMR 2.0, we predict the SMPL pose and shape parameters for each detected unoccluded human [52]. To determine camera translation, we find the intersection between the backprojected ray from the bottom of the 2D bounding box with the ground plane. Given the camera’s intrinsic matrix $\mathbf{K}$, ground plane equation $\begin{bmatrix}\mathbf{n}&d\end{bmatrix}$, and $\mathbf{p}_{b}$ as the pixel coordinate of the bottom of the box, its depth is computed as $z_{b}=-\frac{d}{\mathbf{n}^{T}(\mathbf{K}^{-1}\mathbf{p}_{b})}$. For camera intrinsic parameters and ground plane equation, we employ a recent method by Vuong et al. [73] that utilizes Google Street View [27] to automatically calibrate stationary traffic cameras and infer the ground plane. The mined unoccluded object image with its corresponding segmentation mask, keypoint locations, and 3D poses are used in a clip-art based framework to generate 2D and 3D supervision data in severe occlusions as illustrated in Fig. 2.
@@ 4 Learning Reconstruction under Occlusion
3D Object Reconstruction: To regress object 3D pose and shape under occlusion, we extend the base network with a 3D regression branch. Each vehicle $N$ semantic 3D keypoints $\mathbf{X}$ are represented by a linear combination of the mean shape $\mathbf{\bar{Q}}$ and $K$ principal components $\mathbf{Q}_{1\dots K}$, where $\alpha_{k}$ is the shape coefficients: $\mathbf{X}=\mathbf{\bar{Q}}+\sum_{k=1}^{K}\alpha_{k}\mathbf{Q}_{k}$. Thus, we want to regress for its 6-DoF pose $\{(\mathbf{R},\mathbf{t})\in\mathbf{SE}(3)\}$ and shape coefficients $\alpha_{k}$. Given the predicted 2D keypoints and the corresponding 3D keypoints $\mathbf{X}$, we can obtain its 6-DoF pose through a differentiable PnP layer [9]. Defining $\pi(\cdot)$ as the projection function with known intrinsic $\mathbf{K}$, we can optimize for the keypoint reprojection loss $\mathcal{L}_{\texttt{kp2D}}$, shape loss $\mathcal{L}_{\texttt{shape}}$, and pose loss $\mathcal{L}_{\texttt{pose}}$ as:
@@ 6.3 Ablation Analysis
Cross-evaluations on pedestrian tracking: We evaluate how our WALT3D data helps pedestrian tracking in Table 2. To highlight how better detections lead to better tracking, we use a simple tracker SORT [4] with two detectors: one pretrained on COCO and one finetuned with our WALT3D data. On MOT17-train, the model finetuned with WALT3D data produces high-quality detections under strong occlusions, resulting in improvements over the baseline in all metrics, particularly those favoring better detection quality like MOTA and DetA (see metrics details in [53]).
@@ 7 Discussion and Conclusion
A Summary and More Results
For a brief summary of our method and additional results, we highly encourage the readers to check out the included short video.
B Vehicle 2D Keypoints Dataset
As mentioned in the main paper, although existing datasets like PASCAL3D+ [78], KITTI-3D [47], Carfusion [60], and ApolloCar3D [68] provide annotated 2D vehicle keypoints, they mostly focus on driving scenes [60, 47, 68] or have limited training examples [78], lacking the necessary appearance diversity. To increase the dataset diversity, we prioritized the number of different cameras and viewpoints rather than the number of images per camera. A summary and comparison of our proposed Vehicle 2D Keypoints dataset with other publicly available datasets are detailed in Table 6. On average, we extracted 120 images per camera source for more than 60 different cameras spanning a wide variety of viewpoints, appearances, sensor types, etc. For each image, we run an off-the-shelf object detector to extract the car instances with high confidence score. This set of car instances are manually annotated by the trained annotators from a commercial annotation service. We utilized a web-based interface annotation tool from DeepLabCut [55] where the annotators were asked to select 12 keypoint locations and its corresponding occlusion category (visible/self-occluded/occluded-by-others) for every car. Note that we also asked the annotators to filter out erroneous instances such as bad quality images and/or wrong detections. As of the time of paper submission, we have annotated a total of 42,547 car instances in 7,018 images.
C Camera Intrinsics and Ground Plane
We follow Vuong et al. [73] to obtain the intrinsic parameters and ground plane equation for each of the stationary traffic camera. Specifically, we used the panorama images from Google Street View (GSV) [27] to build a metric 3D scene reconstruction (at the desired camera location), then the stationary camera is registered within the reconstruction to determine its intrinsic and extrinsic parameters. We also geo-register the scene to a metric scale using the GPS coordinates, and the road plane equation is estimated by fitting a plane to the set of 3D points whose 2D pixel projections belong to the road category obtained from off-the-shelf semantic segmentation method [10]. The camera poses and plane equation are used in 3D reconstruction pipeline to reconstruct unoccluded objects as described in the main paper. Thanks to the ground plane geometry constraint, we can reconstruct the accurate 3D geometry of cars and pedestrians, generating realistic occlusion configurations. This method enables us to obtain accurate calibration for more than 100 stationary cameras worldwide, thus allowing for a significant expansion of our clip-art dataset.
D Benchmarking on Additional Datasets
$Acc-\frac{\pi}{6}$ i.i.d shape pose texture context weather NeMo 66.7 51.7 56.9 52.6 51.3 49.8 Ours 75.4 48.6 50.8 56.7 49.1 55.6
Table 3: Comparisons on the OOD-CV [83] dataset (car).
     Refer to caption Refer to caption
Figure 11: Qualitative Results on OccludedPascal3D+ (left) and OOD-CV (right) dataset.
Evaluation on OccludedPascal3D+ dataset: Table 4 shows that our method performs better than NeMo [74] and Ma et al. [54] on the OccludedPascal3D+ [74] dataset.
Method $Acc\ (\frac{\pi}{6})$ $Acc\ (\frac{\pi}{18})$ Med Pose Err Med ADD L1 L2 L3 L1 L2 L3 L1 L2 L3 L1 L2 L3 Occluded PASCAL3D+ Dataset (car) NeMo 48.3 34.3 18.2 17.4 9.6 3.3 0.5 1.0 2.4 1.9 2.0 2.3 Ma et al. 66.6 47.9 27.4 30.8 16.2 5.3 0.3 0.6 1.1 0.8 1.2 1.9 Ours 70.4 56.5 35.3 36.8 25.4 15.3 0.2 0.4 0.8 0.6 1.0 1.4
Table 4: Baseline comparisons across object pose metrics on Occluded-PASDAL3D+ [74] for vehicle category.
Evaluation on OOD-CV dataset: Quantitative results on OOD-CV [83] dataset are shown in Table 3. Although our method has never been trained on the anomalous scenarios in this dataset, our approach shows higher performances on many testing subsets. Please see qualitative results in Fig. 11.
Metric $\delta=0.01$ $\delta=0.1$ $\delta=0.2$ $\delta=0.5$ OC (ours)
Recall 0.60 0.42 0.17 0.01 0.81
Precision 0.32 0.41 0.52 0.57 0.70
Table 5: Accuracy of our OC module compared with baseline using bbox IOU threshold $\delta$ in detecting unoccluded objects.
Mining Unoccluded Objects: To identify unoccluded objects, we evaluate two methods: a simple heuristic based on bounding box IOU threshold $\delta$ (as used in WALT [63]) and training an Occlusion Classifier (OC) using human-annotated data (using images from our new vehicle keypoints dataset). Table 5 demonstrates that our OC module is more effective than the heuristic, particularly in inter-category occlusion scenarios (e.g., vehicles occluded by people or background objects). This allows us to efficiently filter out unwanted occluded objects in the training data, improving data purity. While not essential for our method, we believe this human-annotated dataset is important for future research on understanding and handling occlusion.
Dataset Image source Appearance diversity in terms of # images # car instances Occ. keypoint annotations Per-keypoint occ. type
Cities Times of Day Weathers Viewpoints
PASCAL3D+ Natural Yes Yes Yes No 6,704 7,791 No No
KITTI-3D Self-driving No No No No 2,040 2,040 No No
Carfusion Handheld No No No No 53,000 100,000 Yes No
ApolloCar3D Self-driving No No No No 5,277 60,000 No No
Ours
Handheld
Self-driving
Traffic cameras
Yes Yes Yes Yes 7,018 42,547 Yes Yes
Table 6: Summary and comparison of our Vehicle 2D Keypoints dataset to other publicly available datasets.
E Additional 2D/3D Clip-Art Data Examples
More examples from our 2D/3D Clip-Art pseudo-groundtruth supervision data, including the clip-art image with corresponding amodal segmentation, keypoints, and 3D object reconstruction, are shown in Fig. 12.
F Additional Qualitative Results
Additional results are shown in Fig. 13, with various occlusion configurations, including self-occlusion, truncation, and occlusion-by-others. Notably, training with our clip-art data yields a substantial improvement over baseline methods, particularly in scenarios with heavy occlusion.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Clip-Art Image                                        Segmentation                                   Keypoints                                          3D Reconstruction
Figure 12: Automatically generated 2D and 3D Clip-Art to supervise our network: Unoccluded objects are first mined using time-lapse imagery of the WALT dataset [63]. Random non-intersecting unoccluded objects are composited back into the background image in their respective original positions to preserve correct appearances. The resulting Clip-Art images and their respective amodal segmentation masks, keypoint locations, and 3D meshes are shown. Our method generates realistic appearances from any camera, incorporating diverse viewing geometries, weather conditions, lighting, and occlusion configurations.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Image                Amodal Segmentation     Amodal Keypoints/Shapes                 3D View-1                            3D View-2
Figure 13: We show additional qualitative results on multiple sequences of the WALT [63] dataset. Our method produces accurate amodal segmentation, keypoints, as well as 3D poses and shapes across diverse poses and occlusion configurations.